# Learning to Drive Ethically: Moral Reasoning in Autonomous Driving

### v2

@@ Appendix A Supplementary Material
A.1 Hyperparameter
@@
A.2 Solutions for the longitudinal and lateral polynomial equations
@@
The outputs from the RL agent provide $\dot{l}_{T}$, $d_{T}$, and $T$, while we assume $\ddot{l}_{T}=0$, $\dot{d}_{T}=0$, and $\ddot{d}_{T}=0$. Under these assumptions, the resulting trajectory minimizes jerk, thereby enhancing passenger comfort [Werling et al., 2010]. Therefore, the polynomial coefficients can be solved as follows.
@@ A.3 Scenario analyses
In this scenario, the ego vehicle has right-of-way on the main road, and the right-turning red vehicle must yield. Three baseline agents, standard mode, SACLAG, and EthicAR w/o PER, maintain a relatively high speed through the intersection. After $T=1\mathrm{s}$, their TTC with the red vehicle decreases, driving up the collision risk. By $T=6\mathrm{s}$, both standard mode and EthicAR w/o PER collide with the red vehicle; SACLAG narrowly avoids impact but still operates in a high-risk regime. In contrast, the EthicAR agent behaves more cautiously. At $T=1\mathrm{s}$, it drifts slightly left to create space for the cyclist and settles into a comfortable speed. As the red vehicle initiates its turn, the TTC remains relatively constant, and the risks remain low. Even though the ego vehicle legally had right-of-way, EthicAR proactively yields to the red vehicle, keeping overall risk at a low level.
Figure 8: The real-world scenario in which the ego vehicle (green) travels straight alongside a cyclist (red), while a right-turning vehicle (red) prepares to merge into the main road. (a) are the per-timestep behaviors of EthicAR and other agents, with real-time TTC to the right turn vehicle and maximal risk. (b) is the plot of minimal TTC and maximal risk at the displayed timestep for clear comparison.
@@
Figure 9 depicts a second unprotected left-turn scenario. Here, the ego vehicle (green) is proceeding straight through the intersection, while another vehicle (red) is poised to turn left. Although the ego vehicle nominally has the right of way, an early-arriving or aggressively driven red vehicle may still pose a hazard. How, then, should the ego vehicle modulate its behavior to maintain safety when faced with such unpredictable left-turning traffic?
@@
In standard mode, the ego vehicle fails to slow for oncoming traffic at the intersection, ultimately resulting in a collision. The EthicAR w/o PER agent performs slightly better, its minimal deceleration increases the TTC with the red vehicle, but it still adopts a dangerously aggressive approach. In contrast, both SACLAG and the EthicAR agent demonstrate the strongest safety performance: they proactively reduce the speed as it nears the intersection, thereby minimizing risk to all road users. By yielding courteously and allowing the red vehicle to complete its left turn first.

### v3

@@ Appendix A Supplementary Material
A.1 LSTM-based SACLag Pseudocode
The pseudocode of LSTM-based SACLag with dynamic PER is presented in Algorithm 1.
Algorithm 1 LSTM-based SACLag with Dynamic PER
1:Set hyperparameters: learning rates $\beta_{\theta},\beta_{\phi},\beta_{\lambda},\beta_{\alpha}$, discount $\gamma$, PER exponents $\eta,\beta$, ratio bounds $r_{\min},r_{\max}]$, small constant $\epsilon>0$, soft-update rate $\tau$, and cost threshold $d$.
2:Initialize policy $\pi_{\theta}$, reward critics $Q_{\phi_{1}},Q_{\phi_{2}}$, cost critics $Q_{\psi_{1}},Q_{\psi_{2}}$, target networks $\bar{Q}_{\phi_{1}},\bar{Q}_{\phi_{2}},\bar{Q}_{\psi_{1}},\bar{Q}_{\psi_{2}}$, Lagrange multiplier $\lambda\geq 0$, temperature $\alpha$, PER buffer $\mathcal{D}$.
3:for each iteration do
4:  Sample initial state $s_{0}$ and initial LSTM hidden states $h_{0}^{\pi},h_{0}^{Q}$
5:  for each environment step do
6:   $a_{t}\sim\pi_{\theta}(\cdot|s_{t},h_{t}^{\pi})$,  $s_{t+1}\sim P(\cdot|s_{t},a_{t})$
7:   $\mathcal{D}$ $\leftarrow$ $\mathcal{D}$ $\cup$ $\{s_{t},h_{t}^{\pi},h_{t}^{Q},a_{t},r_{t},c_{t},s_{t+1}\}$ with initial priority
8:  end for
9:  for each gradient step do
10:   Sample batch $\mathcal{B}=\{(s,h^{\pi},h^{Q},a,r,c,s^{\prime})\}$ from $\mathcal{D}$ with PER probabilities $P_{i}\propto p_{i}^{\eta}$
11:   Sample batch of next actions $a^{\prime}\sim\pi_{\theta}(\cdot|s^{\prime},h^{\pi\prime})$
12:   Compute TD errors:
$\delta_{r}=r+\gamma\Big(\min_{j=1,2}\bar{Q}_{\phi_{j}}(s^{\prime},h^{Q\prime},a^{\prime})-\alpha\log\pi_{\theta}(a^{\prime}|s^{\prime},h^{\pi\prime})\Big)-Q_{\phi_{1}}(s,h^{Q},a)$
$\delta_{c}=c+\gamma\min_{j=1,2}\bar{Q}_{\psi_{j}}(s^{\prime},h^{Q\prime},a^{\prime})-Q_{\psi_{1}}(s,h^{Q},a)$
13:   Compute dynamic PER weights:
$r_{e}=\text{clip}\!\left(\tfrac{|\delta_{r}|}{|\delta_{c}|+\epsilon},r_{\min},r_{\max}\right),\;\omega_{r}=\tfrac{r_{e}}{1+r_{e}},\;\omega_{c}=\tfrac{1}{1+r_{e}}$
$p_{i}=\omega_{r}|\delta_{r}|+\omega_{c}|\delta_{c}|,\quad\omega_{i}=\frac{1}{(NP_{i})^{\beta}}$
14:   Update critics with weight $\omega_{i}$
15:   Update actor with Lagrangian:
$\theta\leftarrow\theta-\beta_{\theta}\nabla\Big(w_{i}\,[\,\alpha\log\pi_{\theta}(a|s,h^{\pi})-Q_{\phi_{1}}(s,a)+\lambda Q_{\psi_{1}}(s,a)\,]\Big)$
16:   Update Lagrange multiplier:
$\lambda\leftarrow\max\{0,\,\lambda+\beta_{\lambda}(Q_{\psi_{1}}(s,h^{Q},a)-d)\}$
17:   Update temperature $\alpha$ and target networks with $\tau$
18:   Update PER priorities $p$ in $\mathcal{D}$
19:  end for
20:end for
A.2 Hyperparameter
@@
A.3 Solutions for the longitudinal and lateral polynomial equations
@@
The outputs from the RL agent provide $\dot{l}_{T}$, $d_{T}$, and $T$, while we assume $\ddot{l}_{T}=0$, $\dot{d}_{T}=0$, and $\ddot{d}_{T}=0$. Under these assumptions, the resulting trajectory minimizes jerk, thereby enhancing passenger comfort [werling2010optimal]. Therefore, the polynomial coefficients can be solved as follows.
@@ A.4 Scenario analyses
In this scenario, the ego vehicle has right‑of‑way on the main road, and the right‑turning red vehicle must yield. Three baseline agents, standard mode, SACLAG, and EthicAR w/o PER, maintain a relatively high speed through the intersection. After $T=1\mathrm{s}$, their TTC with the red vehicle decreases, driving up the collision risk. By $T=6\mathrm{s}$, both standard mode and EthicAR w/o PER collide with the red vehicle; SACLAG narrowly avoids impact but still operates in a high‑risk regime. In contrast, the EthicAR agent behaves more cautiously. At $T=1\mathrm{s}$, it drifts slightly left to create space for the cyclist and settles into a comfortable speed. As the red vehicle initiates its turn, the TTC remains relatively constant, and the risks remain low. Even though the ego vehicle legally had right‑of‑way, EthicAR proactively yields to the red vehicle, keeping overall risk at a low level.
Figure 9: The real-world scenario in which the ego vehicle (green) travels straight alongside a cyclist (red), while a right‑turning vehicle (red) prepares to merge into the main road. (a) are the per-timestep behaviors of EthicAR and other agents, with real-time TTC to the right turn vehicle and maximal risk. (b) is the plot of minimal TTC and maximal risk at the displayed timestep for clear comparison.
@@
Figure 10 depicts a second unprotected left‑turn scenario. Here, the ego vehicle (green) is proceeding straight through the intersection, while another vehicle (red) is poised to turn left. Although the ego vehicle nominally has the right of way, an early‑arriving or aggressively driven red vehicle may still pose a hazard. How, then, should the ego vehicle modulate its behavior to maintain safety when faced with such unpredictable left‑turning traffic?
@@
In standard mode, the ego vehicle fails to slow for oncoming traffic at the intersection, ultimately resulting in a collision. The EthicAR w/o PER agent performs slightly better, its minimal deceleration increases the TTC with the red vehicle, but it still adopts a dangerously aggressive approach. In contrast, both SACLAG and the EthicAR agent show the strongest safety performance: they proactively reduce the speed as it nears the intersection, thereby minimizing risk to all road users. By yielding courteously and allowing the red vehicle to complete its left turn first.